OpenStudy (anonymous):

Simplify using your knowledge of properties. 2(a+4)+ -2a+-8

OpenStudy (unklerhaukus):

\[2(a+4)+ -2a+-8\] \[2(a+4) -2(a+4)\] \[(2-2)(a+4)\]

OpenStudy (anonymous):

what is the final answer, I get -4

OpenStudy (unklerhaukus):

HOW?

OpenStudy (unklerhaukus):

2-2=0 \[(2−2)(a+4)=(0)(a+4)=0\times(a+4)=\cdots?\]

OpenStudy (anonymous):

so, 2-2=0 and 0xa+4=0 also

OpenStudy (unklerhaukus):

yer

OpenStudy (anonymous):

ok, got it, thanks